OpenStudy (anonymous):

In a right triangle the legs are 10 and 24. Find the length of the hypotenuse.

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2 \) \(10^2 + 24^2 = c^2 \)

OpenStudy (anonymous):

I got 26?

OpenStudy (mathstudent55):

Correct. The hypotenuse has length 26.